WATER’S ROLE IN RESHAPING A MACROCYCLE’S BINDING POCKET: CONFORMATION-SPECIFIC INFRARED AND ULTRAVIOLET SPECTROSCOPY OF BENZO-15-CROWN-5-(H_{2}O)_{n}-CLUSTERS (n = 1, 2)

V. ALVIN SHUBERT, Argonne National Laboratory, Chemical Sciences and Engineering Division, 9700 South Cass Avenue, Argonne, IL 60439; \underline{CHRISTIAN W. MÜLLER}, WILLIAM H. JAMES III and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-2084.

Crown ethers are well-studied examples of flexible macrocycles with a high binding selectivity for substrates, especially cations. We investigated the conformational preferences of the singly and doubly complexed water clusters of the crown ethers benzo-15-crown-5 (B15C) and its amino-derivative 4’-aminobenzo-15-crown-5 (ABC) cooled in a supersonic jet expansion. The fluorescence excitation, resonance enhanced two-photon ionization (R2PI), UV-UV holeburning (UVHB), fluorescence-dip infrared (FDIR), resonant ion-dip infrared (RIDIR) and novel IR-IR-UV holeburning^{1} spectra allowed for the identification of two B15C–(H_{2}O)_{1} conformers and one ABC–(H_{2}O)_{1} conformer. These conformers are characterized by an all-planar arrangement of the atoms directly bound to the benzene ring in which the crown ether macrocycle opens up to a symmetric structure and accomodates a doubly and triply H-bonded H_{2}O molecule in two distinct ways, respectively. Two B15C–(H_{2}O)_{2} conformers and one ABC– (H_{2}O)_{2} conformer were identified. One of the B15C–(H_{2}O)_{2} conformers contains a macrocycle configuration identical to that found in the monohydrated clusters with an H-bonding topology in which the H_{2}O molecules occupy both available sites simultaneously. The second B15C–(H_{2}O)_{2} conformer is assigned to an H-bond pattern in which the two H_{2}O molecules are concatenated to form an H-bonded bridge involving only three of the four available O–Hbonds (see figure).

(1) V. A. Shubert and T. S. Zwier, J. Phys. Chem. A, 2007, 111, 13283.